### v3

# SystolicAttention: Fusing FlashAttention within a Single Systolic Array

Abstract.
Transformer models rely heavily on scaled dot-product attention (SDPA), typically implemented using the FlashAttention algorithm. However, current systolic-array-based accelerators face significant challenges when executing FlashAttention. Systolic arrays achieve high utilization primarily for consecutive and large matrix multiplications, whereas FlashAttention requires frequent interleaving of matrix multiplications and softmax operations.
The frequent data swaps between matrix multiplications on the systolic array and softmax operations on external units result in low array utilization. Moreover, when these computations run concurrently, the softmax stage contends with matrix multiplication for register file and SRAM ports, further degrading performance.
To overcome these limitations, we propose FSA, an enhanced systolic array architecture that enables the FlashAttention algorithm to run entirely within a single systolic array—eliminating the need for external vector units. At the core of FSA is SystolicAttention, a novel scheduling algorithm that maps FlashAttention operations onto systolic arrays with fine-grained, element-wise overlap. This approach significantly improves array utilization while preserving the original floating-point operation order to maintain numerical stability.
[…]
Systolic Array, Flash Attention, Microarchitecture, Accelerators
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computer systems organization Systolic arraysccs: Hardware
1. Introduction
The rapid advancement of artificial intelligence, particularly in natural language processing and computer vision, has been largely driven by the success of transformer models (Vaswani et al., 2017; Devlin et al., 2019; Brown et al., 2020), which have revolutionized applications ranging from language translation and text generation (Ouyang et al., 2022) to image recognition (Kolesnikov et al., 2021) and autonomous driving (Carion et al., 2020). However, the computational demands of transformer models are enormous (Strubell et al., 2019), requiring specialized hardware accelerators to achieve practical performance and energy efficiency. As these models continue to grow in size and complexity (Kaplan et al., 2020), the need for efficient hardware acceleration becomes increasingly critical. One of the key operations in transformer models is the scaled dot-product attention (SDPA), which is often implemented using the FlashAttention (Dao, 2023; Shah et al., 2024) algorithm. FlashAttention reduces memory overhead by breaking the $Q$, $K$, and $V$ matrices into small tiles and fusing matrix multiplication with online softmax computation.
Figure 1. Active time percentage of different components in AWS NeuronCore-v2 when running FlashAttention.
[…]
Active Time
Systolic arrays (Kung, 1982) are specialized hardware architectures that perform computation by streaming data through a network of tightly coupled processing elements (PEs). Many systolic array based accelerators with various dataflows (Chen et al., 2017, 2016; Kao et al., 2023; Nayak et al., 2024) have been proposed to accelerate deep learning workloads. In this paper, we focus on multiply-accumulate (MAC) based 2D systolic arrays, which are the most widely used in commercial AI accelerators, such as Google’s TPU (Google Cloud, 2024; Jouppi et al., 2017, 2023) and AWS NeuronCore (AWS Neuron Documentation, 2024b). In these architectures, data moves only one hop per clock cycle, minimizing hardware data movement overhead and enabling high scalability, which is crucial for efficiently executing large matrix multiplications.
When mapping FlashAttention to systolic array based accelerators, the matrix multiplication operations can be executed using the systolic array, while softmax-related operations must be offloaded to external vector and scalar units. To maximize hardware utilization, matrix multiplication and softmax operations from different iterations are often software-pipelined to run in parallel. Thus, the vector and scalar units must provide sufficient floating-point operations per second (FLOPs/s) to keep up with the systolic array. However, non-matrix-multiplication FLOPs/s are typically much lower in volume and more expensive to improve in hardware, resulting in a significant performance bottleneck. As shown in Figure 1, when running FlashAttention on AWS NeuronCore-v2, the systolic array (tensor engine) is only active for about 45% of the time, while the scalar unit remains active for 80% of the time.
Beyond the imbalance between matrix and non-matrix FLOPs/s, the complex data dependencies in FlashAttention also contribute to low systolic array utilization. Systolic arrays achieve high utilization when matrix data can be reused across multiple iterations, improving data locality and reducing the preloading overhead. However, in FlashAttention, the two matrix multiplications are separated by the softmax operation, preventing reuse of intermediate data. The softmax operation also causes frequent data transfers between the systolic array and the vector unit, further increasing preloading overhead and SRAM port contention (Kim et al., 2025; Xu et al., 2025). As shown in our experiments in subsection 6.1, the systolic array achieves less than 25% FLOPs/s utilization, even with 45% active time on AWS NeuronCore-v2, significantly lower than the theoretical peak FLOPs/s of the array.
The above problems can be addressed by executing the full FlashAttention algorithm entirely on the systolic array. In this case, non-matrix-multiplication operations can be computed using the matrix multiplication FLOPs/s available on the array itself. Additionally, data preloading overhead and SRAM port contention are eliminated since all computations occur within the same array. Achieving this, however, requires overcoming two key challenges. First, non-matrix-multiplication operations such as exp, rowmax, and rowsum must be efficiently implemented using only the systolic array. Second, different operations must be overlapped effectively to maximize utilization of the array.
COSA Plus (Wang et al., 2025) proposes mapping FlashAttention to two cooperative systolic arrays, but still requires the softmax operation to be performed by a special function unit (SFU), which may not provide sufficient FLOPs/s to keep up with the systolic arrays. FuseMax (Nayak et al., 2024) pioneered the fusion of FlashAttention into a single systolic array. It maps the rowmax and rowsum operations as spatial reductions and overlaps multiple FlashAttention iterations on the array simultaneously. However, it requires storing intermediate results using 10 registers per PE to support context switching between iterations, leading to additional register overhead.
We propose a new hardware-friendly systolic array architecture, FSA, which introduces minimal overhead to existing commercial systolic arrays and enables the execution of the full FlashAttention algorithm on a single systolic array. This design is guided by the following key insights: (1) the input values to the exp operation in FlashAttention are always less than or equal to zero. By splitting the input into integer and fractional parts, the fractional component falls within the range (1,0](-1,0], making it suitable for approximation via linear interpolation (Koca et al., 2023) reusing the MACs available on the systolic array; (2) the rowmax and rowsum operations can be expressed as reduction operations, which are well-supported in systolic architectures and can be performed on-the-fly as rows are generated in a streaming fashion; and (3) the operations within each FlashAttention iteration exhibit a dataflow pattern that allows fine-grained execution overlap, enabling multiple computations within the same iteration to proceed in parallel without waiting for the completion of prior steps. We also introduce SystolicAttention, a scheduling algorithm that makes the best use of FSA’s capabilities to overlap element-wise operations, significantly improving array utilization.
Our key contributions are as follows:
  • We present FSA, an enhanced systolic array architecture capable of executing the full FlashAttention algorithm without relying on additional vector or scalar units. FSA is implemented in synthesizable RTL, requiring only 12% additional area compared to a standard design. The RTL implementation achieves a clock frequency of 1.5 GHz on a 16 nm process and 70 MHz on the Xilinx Alveo U55C FPGA.
[…]
  • We develop a Python programming interface that enables users to write custom FSA kernels. Both the RTL implementation and the Python software stack are open-sourced at https://github.com/VCA-EPFL/FSA.
2. Background and Motivation
In this section, we first briefly introduce baseline systolic arrays and how they perform matrix multiplication. We then present the FlashAttention algorithm and discuss the challenges of implementing it on systolic array based accelerators. Finally, we elaborate on the insights and challenges of running the full FlashAttention algorithm entirely within a systolic array.
2.1. Systolic Array Architectures
Systolic arrays are a class of hardware accelerators that perform computations in a pipelined manner, where data flows through an array of processing elements (PEs) in a rhythmic fashion. Each PE performs a simple operation, such as a multiply-accumulate (MAC), and the data flows through the array in a wave-like pattern. Systolic arrays are particularly well-suited for matrix operations, such as matrix multiplication, as they can efficiently exploit data parallelism and locality. The dataflow of a systolic array defines how data moves through the array and the communication pattern between PEs. Common dataflows include weight-stationary, input-stationary, output-stationary, and row-stationary (Chen et al., 2016). Due to the transpose rule of matrix products, weight-stationary and input-stationary dataflows can often be implemented using the same hardware. In this paper, we use the term weight-stationary to refer to both. As most deep learning workloads favor weight reuse, many systolic accelerators adopt the weight-stationary dataflow, such as Google’s TPU (Google Cloud, 2024).
2.2. Matrix Multiplication on Systolic Arrays
An example weight-stationary systolic array is shown in Figure 2. It consists of $N\times N$ PEs, where each PE performs one floating-point MAC per cycle. One input matrix is tiled into $N\times N$ and preloaded into the array (stationary matrix). The other is tiled into $N\times M$ and streamed in from SRAM (moving matrix). The output is similarly tiled into $N\times M$ and streamed out.
Figure 2. Computing $C\mathrel{+}=AB$ on a weight-stationary systolic array.
[…]
SA
Since the weight-stationary array does not retain the output matrix internally, output tiles are stored in a small, dedicated accumulation SRAM located at the bottom of the array. An accumulator unit reads values from this accumulation SRAM, combines the newly computed results from the systolic array with existing values, and writes the updated results back. This near-memory accumulation scheme is widely adopted in both open-source accelerators such as Gemmini (Genc et al., 2022), and commercial products such as AWS NeuronCore (AWS Neuron Documentation, 2024b).
Preloading the stationary matrix requires $N$ cycles at the start of computation. To ensure correctness, each row of the moving matrix must be delayed by 0 to N1N-1 cycles before entering the array, and each column of the output matrix must be delayed by 0 to N1N-1 cycles before being read out. This introduces a total synchronization overhead of 2N12N-1 cycles. The total computation takes M+3N1M+3N-1 cycles, resulting in an array utilization of $\frac{M}{M+3N-1}$. When $M\gg N$, this preloading and synchronization overhead is effectively amortized over the duration of the computation.
2.3. Implementing FlashAttention on Systolic Array Based Accelerators
Require: Matrices $Q,K,V\in\mathbb{R}^{LEN\times d}$
1 Divide $Q$ into $T_{r}=\lceil\frac{LEN}{B_{r}}\rceil$ blocks of size $B_{r}\times d$ each, and divide $K$ and $V$ into $T_{c}=\lceil\frac{LEN}{B_{c}}\rceil$ blocks of size $B_{c}\times d$ each;
2 for $1\leq i\leq T_{r}$ do
3   Initialize $old\_m,old\_l=(-\infty),(0)\in\mathbb{R}^{B_{r}}$;
4   Initialize $old\_O=(0)\in\mathbb{R}^{B_{r}\times d}$;
5 for $1\leq j\leq T_{c}$ do
6    $S=Q_{i}K_{j}^{T}\in\mathbb{R}^{B_{r}\times B_{c}}$ ;
7    $local\_m=rowmax(S)\in\mathbb{R}^{B_{r}}$;
8    $new\_m=max(local\_m,old\_m)\in\mathbb{R}^{B_{r}}$;
9    $a=old\_m-new\_m\in\mathbb{R}^{B_{r}}$;
10    $b=exp(\frac{a}{\sqrt{d}})=exp2(\frac{\log_{2}e}{\sqrt{d}}a)\in\mathbb{R}^{B_{r}}$;
11    $N=S-new\_m\in\mathbb{R}^{B_{r}\times B_{c}}$;
12    $P=exp(\frac{N}{\sqrt{d}})=exp2(\frac{\log_{2}e}{\sqrt{d}}N)\in\mathbb{R}^{B_{r}\times B_{c}}$;
13    $local\_l=rowsum(P)\in\mathbb{R}^{B_{r}}$;
14    $new\_l=old\_l\times b+local\_l\in\mathbb{R}^{B_{r}}$ ;
15    $local\_O=PV_{j}\in\mathbb{R}^{B_{r}\times d}$ ;
16    $new\_O=diag(b)old\_O+local\_O\in\mathbb{R}^{B_{r}\times d}$;
17    old_m=new_mold\_m=new\_m;
18    old_l=new_lold\_l=new\_l;
19    old_O=new_Oold\_O=new\_O;
20
21   end for
22 $O_{i}=diag(old\_l)^{-1}old\_O\in\mathbb{R}^{B_{r}\times d}$;
23
24 end for
Algorithm 1 FlashAttention-2 and 3 forward pass
The FlashAttention algorithm (Dao, 2023; Shah et al., 2024) is shown in algorithm 1, focusing on the computational steps and omitting memory load/store operations for brevity. Matrix multiplications are highlighted in red, while non-matrix operations such as reductions and element wise computations are shown in blue. The exponential function is implemented as $\exp(x)=2^{x\log_{2}e}$, following the official open-source FlashAttention implementation, since exp2 is typically more efficient than exp on hardware accelerators.
To implement FlashAttention on systolic array based accelerators, the first matrix multiplication result must be moved out to a vector unit or scalar unit for softmax-related operations. Intermediate results must then be sent back into the array for the second matrix multiplication. As discussed in subsection 2.2, these round-trips incur preloading and synchronization overheads. To amortize them, large tile sizes are preferred. However, larger tiles increase the imbalance between matrix and non-matrix FLOPs and raise SRAM demand for intermediate storage. Even with software pipelining to overlap execution, the FLOPs/s imbalance between matrix multiplications and other operations leads to substantial time spent on non-matrix computations (Nayak et al., 2024). Achieving high performance for FlashAttention on systolic arrays remains challenging.
2.4. Motivation
Running the complete FlashAttention algorithm within a single systolic array addresses several critical inefficiencies. By eliminating offloading to external vector or scalar units, all computations can leverage the abundant matrix multiplication FLOPs/s inherently available within the systolic array, significantly improving overall hardware utilization. However, achieving full execution of FlashAttention within the array presents considerable challenges:
  • Efficient Non-MAC Operations: Operations such as exp, rowmax, and rowsum typically rely on external units due to their element-wise or reduction nature. Implementing these efficiently using only MAC-based PEs requires careful algorithmic transformations and architectural extensions.
  • Computation Overlap: To maximize utilization, non-matrix operations must be tightly overlapped with ongoing matrix computations. The complex data dependencies within FlashAttention demand a finely tuned scheduling strategy to ensure that overlapping computations do not stall each other.
  • Dataflow and SRAM Management: The intermediate data produced by the first matrix multiplication in FlashAttention must be effectively reused without excessive SRAM accesses or port contention. Achieving this reuse within the constrained internal storage of the systolic array poses a significant design challenge.
Addressing these challenges requires novel architectural enhancements and sophisticated scheduling algorithms, forming the basis for our proposed FSA architecture and SystolicAttention scheduling algorithm detailed in subsequent sections.
3. FSA Overview
The design principle of FSA is to enable the execution of the full FlashAttention algorithm on a single systolic array with minimal hardware overhead, thereby improving the FLOPs/s for the non-matrix multiplication operations and eliminating the need for an external vector unit. To achieve this, we add a row of comparison units at the top of the array to perform the on-the-fly rowmax operation. Additionally, we introduce a Split unit in each PE to implement the exp operation. To allow the rowmax, exp, and rowsum operations to overlap with matrix multiplication, we incorporate an upward data path in the systolic array.
In this section, we first provide an overview of the enhanced systolic array microarchitecture, FSA, that we propose to accelerate the FlashAttention algorithm. We then describe how the non-matrix operations in FlashAttention can be executed using FSA. Finally, we demonstrate how to extensively overlap different computations within FSA in a systolic manner using our proposed SystolicAttention scheduling algorithm.
3.1. FSA Design Overview
Figure 3. FSA’s modifications to standard systolic array.
[…]
FSA’s modifications.
As shown in Figure 3, FSA is a systolic array based accelerator that extends the standard architecture with three key modifications to enable efficient execution of FlashAttention:
  1. (1)
    A row of comparison units at the top of the array to compute the column-wise maximum of the output on the fly.
  2. (2)
    A Split unit in each PE to decompose the input floating-point value into its integer and fractional parts.
  3. (3)
    An upward data path to propagate partial sums in the reverse direction.
The detailed design of these modifications is as follows:
Row of comparison units: To implement lines 11 of algorithm 1, we add a row of comparison units at the top of the array to compute the maximum value of each output column dynamically. Each unit consists of a floating-point adder and two registers. The register old_mold\_m stores the maximum value from the previous iteration, while new_mnew\_m holds the temporary maximum of the current iteration. The value of new_mnew\_m is updated by comparing old_mold\_m and the current input using subtraction through the floating-point adder. This differential computation enables efficient max tracking inline with systolic flow.
Split unit in each PE: Each PE is equipped with a Split unit that separates the input floating-point number into its integer and fractional components. The exp2 operation is applied to the fractional part using piecewise linear interpolation, which is efficiently handled by the PE’s MAC unit.
Upward data path: Standard systolic arrays propagate partial sums in one direction only. We add an additional data path that enables upward propagation, allowing computation to start from either the top or bottom of the array.
These architectural enhancements allow FSA to support operations that are not feasible on a baseline systolic array, thereby enabling efficient standalone execution of FlashAttention, as elaborated in the following sections.
3.2. rowmax on FSA
Figure 4. On-the-fly rowmax generation.
[…]
On-the-fly rowmax generation
Figure 4 illustrates the fused matrix multiplication and rowmax computation, which corresponds to lines 11 of algorithm 1. The matrix multiplication $S=Q_{i}K^{T}_{j}$ is performed in the systolic array along the upward data path, such that the rows of $S$ become the columns of the systolic array and exit the array from the top.
The rowmax operation is executed in the comparison units concurrently with the generation of each row of $S$. As the rows of $S$ are produced, they are re-streamed from the top of the array back down through the comparison units. By the time the maximum value of each row is computed, $S$ is already resident in the systolic array.
Consequently, the subtraction operation N=Snew_mN=S-new\_m from algorithm 1 of algorithm 1 can be performed immediately. This is achieved by streaming constant ones from the left of the array as the multiplicand, and streaming new_mnew\_m from the top of the array as the addend.
3.3. exp and exp2 on FSA
After the rowmax and subtraction operations, an element-wise exponential function must be applied to the matrix $N$, as shown in algorithm 1 of algorithm 1:
$P=exp\left(\frac{N}{\sqrt{d}}\right)=exp2\left(\frac{\log_{2}e}{\sqrt{d}}\cdot N\right).$
The exp function is implemented using exp2, which is typically more efficient on hardware accelerators (Hussain et al., 2022). The constant multiplication $\frac{\log_{2}e}{\sqrt{d}}\cdot N$ can be computed by streaming the constant $\frac{\log_{2}e}{\sqrt{d}}$ from the left of the array as the multiplicand, since the matrix $N$ is already resident in the systolic array after the subtraction operation. For the element-wise exp2 operation, we use a piecewise linear interpolation to approximate the function (Strollo et al., 2011).
As proposed in (Koca et al., 2023), the input to exp2 can be split into its integer and fractional parts for a fixed-point number, and the fractional part can be approximated using piecewise linear interpolation. We found that this approach is also applicable to floating-point numbers. The piecewise linear interpolation can be efficiently implemented by reusing the MAC units in the systolic array. A floating-point number $x$ can be decomposed into its integer and fractional components: $x=x_{i}+x_{f}$, where $x_{i}$ is the integer part and $x_{f}$ is the fractional part (iee, 2019). This decomposition leads to the identity:
$exp2(x)=2^{x_{i}+x_{f}}=2^{x_{i}}\cdot 2^{x_{f}}.$
Since $N=S-\texttt{rowmax}(S)$, all elements in $N$ are less than or equal to zero. This implies $x_{f}\in(-1,0]$, and therefore $2^{x_{f}}\in(0.5,1]$. Given this limited range for $x_{f}$, we find that an 8-piece uniform piecewise linear interpolation is sufficient to achieve a relative error around $10^{-2}$ in the final FlashAttention results:
$exp2(x)\approx 2^{x_{i}}\cdot\left(slope_{k}\cdot x_{f}+intercept_{k}\right),\quad k\in[0,8).$
As shown in Figure 5, the piecewise linear interpolation for $x_{f}$ can be implemented by reusing the MAC units in the systolic array. A simple Split unit is added to separate the input $x$ into its integer and fractional parts. The Split unit extracts $x_{i}$ and $x_{f}$ by shifting the mantissa bits to align the exponent to zero. The term $2^{x_{i}}$ affects only the exponent bits in the final result, effectively increasing the exponent by $x_{i}$.
Figure 5. exp2 using piecewise linear interpolation.
[…]
exp2 using piecewise linear interpolation.
To avoid the need for storing the piecewise linear interpolation coefficients in the systolic array, we stream the $slope_{k}$ and $intercept_{k}$ values from the left and top of the array, respectively. We further observe that all intercepts lie in the range $(0.5,1]$, so their exponent can only be 0 or 1-1. The MSBs of their exponents can be used to encode the interpolation index $k$. This enables each PE to update its register based on $k$ without requiring additional control signals.
3.4. rowsum on FSA
Benefiting from the in-place exp computation of $P=exp\left(\frac{N}{\sqrt{d}}\right)$ in the systolic array, the rowsum operation can begin as soon as the top-left element of $P$ is produced.
Figure 6. rowsum using systolic array.
[…]
rowsum using systolic array.
As shown in Figure 6, the rowsum of $P$ can be computed by streaming constant ones from the left of the array as the multiplicand, and streaming zeros from the top of the array as the addend. The second matrix multiplication, $O=PV$ can then be performed in the systolic array along the downward data path, starting one cycle after the rowsum operation begins. Once the rowsum local_llocal\_l and the output local_Olocal\_O exit the array from the bottom, they are accumulated with the previous values old_lold\_l and old_Oold\_O in the accumulator, and the results update new_lnew\_l and new_Onew\_O in the accumulation SRAM, respectively.
3.5. Overlapping Computations with SystolicAttention
Figure 7. Overlapped computations of the FlashAttention inner loop in the systolic array.
[…]
Overlapping computations in the systolic array.
So far, we have introduced the key operations required to perform FlashAttention within a systolic array, including rowmax, exp2, and rowsum. We now describe how these operations can be overlapped in a systolic fashion to maximize efficiency.
Each iteration of the FlashAttention algorithm involves two matrix multiplications. To avoid costly data movement between the systolic array and external SRAM, we structure computation on an $N\_ROWS\times N\_COLS$ systolic array by selecting tile sizes $B_{r}=N\_COLS$ and $B_{c}=N\_ROWS=d$. This allows the output of the first matrix multiplication, $S=Q_{i}K_{j}^{T}$, to be directly re-streamed into the array as input for the second matrix multiplication.
During the first matrix multiplication, the head dimension $d$ of both $Q$ and $K$ is mapped to N_ROWSN\_ROWS, while the sequence length $B_{r}$ of $Q$ is mapped to N_COLSN\_COLS. After computing $S$, we re-stream it from the top of the array, mapping the $B_{c}$ dimension of $S$ to N_ROWSN\_ROWS. A sequence of element-wise operations is then performed in-place cycle by cycle to transform $S$ into $N$ and then into $P$.
For the second matrix multiplication, $O=PV$, the $B_{c}$ dimension of $V$ is mapped to N_ROWSN\_ROWS and contracted with the stationary matrix $P$ inside the systolic array.
We use the upward data path for the first matrix multiplication, such that the elements of $P$ are generated from the top left to the bottom right of the array. Then, by leveraging the downward data path, we can immediately begin the second matrix multiplication once the top-left element of $P$ is available.
Figure 7 illustrates the execution of the entire inner loop. The fifth operation, a=old_mnew_ma=old\_m-new\_m, passes down the difference between the previous and current maximum values from comparison units to the accumulator. This operation runs in parallel with the constant multiplication $\frac{\log_{2}e}{\sqrt{d}}\cdot N$, as the downward data path is idle during the in-place constant multiplication.
Overall, each iteration takes only $2\times N\_COLS+3\times N\_ROWS+10$ cycles, meaning that on an $N\times N$ systolic array, it takes just $5N+10$ cycles to process a $N\times N$ FlashAttention tile. By contrast, on a naive $N\times N$ systolic array, the two independent matrix multiplications alone may require up to 8N28N-2 cycles due to preloading and synchronization overheads discussed in subsection 2.2.
The re-scaling operation in algorithm 1 of algorithm 1 is only performed once per outer loop. It is executed using the accumulator after the second matrix multiplication completes. In our implementation, this re-scaling step takes $2N+20$ cycles, which is negligible compared to the inner loop execution time.
4. FSA Microarchitecture
FSA modifies the standard systolic array architecture to support the SystolicAttention algorithm. To validate its feasibility and quantify the associated hardware overhead, we implement FSA in RTL and design a minimal yet flexible instruction set capable of executing the full FlashAttention.
The RTL design is implemented in Chisel HDL (Bachrach et al., 2012) and built using the Chipyard (Amid et al., 2020b, a, 2021; Cook et al., 2017) framework for system-on-chip (SoC) construction. Unlike prior efforts such as Gemmini (Genc et al., 2022), which focus on flexible design space exploration, our goal is to deliver a concrete and complete implementation of FSA to validate both the practicality of the SystolicAttention algorithm and its hardware cost.
This section provides an overview of the FSA microarchitecture, details the instruction set design that simplifies the hardware implementation, and describes the systolic array controller that orchestrates the execution of the SystolicAttention algorithm.
4.1. Microarchitecture Overview
Figure 8. FSA microarchitecture.
[…]
FSA microarchitecture overview.
An overview of the FSA microarchitecture is shown in Figure 8. FSA receives instructions via an AXI4-Lite interface and accesses backing memory through a configurable number of AXI4 memory interfaces.
The instruction decoder classifies instructions into three types: load, store, and compute. Instructions of different classes are executed asynchronously, while instructions of the same class are issued in order but their execution may be interleaved.
The systolic array controller issues compute instructions once the required data has been loaded into SRAM. To simplify control logic, FSA prioritizes SRAM access over compute operations, making the latency of compute instructions fully deterministic once issued. This design allows the controller to statically schedule control signals based on the known execution cycle of each instruction.
The DMA engine follows an iDMA-style interface (Benz et al., 4 01) and supports 2D data transfers. Each DMA instruction can specify source and destination addresses in SRAM and backing memory, along with the transfer size and stride.
When multiple AXI4 channels are available, the DMA engine automatically partitions the transfer into parallel AXI4 transactions, maximizing memory bandwidth utilization.
4.2. FSA Instruction Set
To enable fine-grained overlap between various FlashAttention operations, the systolic array in FSA requires more sophisticated control logic than standard systolic arrays. The primary design goal of the FSA instruction set is to ensure that compute instructions execute in a fully deterministic manner, independent of memory access latency. This simplification greatly reduces control complexity.
Figure 9. FSA compute instructions.
[…]
FSA compute instructions overview.
To achieve this, we define five compute instructions based on their SRAM data dependencies, as illustrated in Figure 9. The FlashAttention inner loop is divided into three phases:
  • Load Stationary: Preload $Q$ into the systolic array.
  • Attn Score: Perform the first matrix multiplication using $Q$ and $K$, fused with element-wise online softmax to compute the $P$ matrix in-place within the systolic array. The log of the exponent sum for $P$ is also calculated during this step.
  • Attn Value: Perform the second matrix multiplication using $V$ and $P$.
The FlashAttention outer loop is divided into two additional phases:
  • Reciprocal: Load the log exponent sum of $P$ from the accumulation SRAM and compute its reciprocal. The result is stored in the accumulator as a scaling factor for the next phase.
  • Attn LSE Norm: Load the output matrix $O$ from the accumulator and apply the reciprocal scaling factor.
Each compute instruction reads one input tile from SRAM and writes one output tile to the accumulator SRAM. This one-tile-in, one-tile-out design ensures that compute instructions can be executed as soon as their corresponding SRAM tile is ready—without waiting on other tiles to be loaded or processed.
DMA instructions follow a similar structure to compute instructions, consisting of one memory tile descriptor and one SRAM tile descriptor. However, they use wider bit fields to support large memory address spaces.
4.3. Systolic Array Controller
Figure 10. FSA systolic array controller.
[…]
FSA controller overview.
The microarchitecture of the FSA systolic array controller is illustrated in Figure 10. The controller comprises two identical counter-based finite state machines (FSMs), each responsible for generating control signals for a compute instruction.
Control signals are categorized into three types:
  • PE Control Signals: Define the operations to be performed within the PEs of the systolic array.
  • Accumulator Control Signals: Specify the actions performed in the accumulator.
  • SRAM Control Signals: Govern the read and write operations to SRAM required for executing the compute instruction.
Because compute instruction execution is fully deterministic, the counter is initialized at the start of each compute instruction, incremented every cycle, and reset to zero upon completion. As a result, the FSMs can generate control signals solely based on the instruction type and the current counter value.
To simplify control logic development, we implemented a domain specific language (DSL) on top of Chisel and Scala. This DSL allows designers to specify which control signals should be scheduled at each cycle. The control logic is then automatically synthesized from the DSL specification.
The dual-FSM architecture enables overlapping execution of two compute instructions. For example, the Attn Value instruction can begin as soon as the Attn Score instruction produces the first element of the $P$ matrix. A combiner unit merges the control signals from both FSMs, ensuring that signals are emitted in the correct order and that conflicts are avoided.
5. FSA Kernel Programming Model
The AWS Neuron Kernel Interface (NKI) (AWS Neuron Team, 2025) provides an expressive Python interface that allows users to write custom kernels for AWS Neuron devices. It offers instruction-level control with the convenience of Python APIs.
Inspired by AWS NKI, FSA provides a similar, simplified Python library for writing custom FSA kernels. The library consists of three main components: type-safe tensors, a Python API for the FSA instruction set, and a lightweight JIT compiler.
5.1. Type-Safe Tensors
The FSA Python library implements a PyTorch-like (Paszke et al., 2019; Ansel et al., 2024) tensor abstraction for representing multi-dimensional arrays. Because the memory hierarchy is fully user-managed, tensors can be allocated in one of three memory spaces: main memory, scratchpad SRAM, or accumulation SRAM. Accordingly, we define three tensor types:
  • MTile — for main memory tensors,
  • STile — for scratchpad SRAM tensors,
  • ATile — for accumulation SRAM tensors.
All tensor types support a subset of the PyTorch API, including shape, dtype, split, and reverse. By explicitly distinguishing tensor types, users can write custom kernel functions that clearly declare the expected input and output tensor memory scopes.
5.2. Python API for FSA Instructions
Each FSA instruction is exposed through a corresponding Python API. These APIs are designed to be type-safe, ensuring correct usage of tensor types. The complete set of supported APIs is shown in 1.
Listing 1: FSA Python instructions.
# DMA Instructions
def store_tile(src: ATile, dst: MTile, ...)
def load_tile(src: MTile, dst: STile, ...)
# Compute Instructions
def load_stationary(tile: STile, ...)
def attn_score(K: STile, l: ATile, ...)
def attn_value(V: STile, O: ATile, ...)
def reciprocal(l: ATile, ...)
def attn_lse_norm(O: ATile, ...)
Internally, these functions construct hardware instructions using the metadata of the input tensors—such as shape, stride, and data type. This abstraction allows users to write high-level kernels without managing low-level hardware details.
5.3. JIT Kernel Compiler
We provide a lightweight JIT compiler that compiles decorated Python functions into FSA binary programs. These programs consist of a sequence of binary instructions that can be submitted to the device’s instruction queue for execution.
The FlashAttention kernel implemented in FSA using the JIT compiler is shown in 2.
Listing 2: FSA FlashAttention kernel.
import fsa as F
@F.kernel(device="verilator_sim")
def attention(Q: MTile, K: MTile, Vt: MTile) -> MTile:
# allocate output tensor
Ot: MTile = F.alloc_mem(..)
# split large tensors into tiles
Ot_MTiles = Ot.split(br, dim=-1) # [d, br]
Q_MTiles = Q.split(br, dim=-2) # [br, d]
K_MTiles = K.split(bc, dim=-2) # [br, d]
Vt_MTiles = Vt.split(bc, dim=-1) # [d, bc]
# double buffering for Q, K, Vt
K_STiles = (F.alloc_spad(..), F.alloc_spad(..))
Vt_STiles = (F.alloc_spad(..), F.alloc_spad(..))
# allocate space for accumulation results
log_expsum = F.alloc_accum(..) # [1, br]
Ot_ATile = F.alloc_accum(..) # [d, br]
for i, Q_i in enumerate(Q_MTiles):
F.load_tile(Q_i, Q_STiles[i % 2])
for j, (K_j, Vt_j) in enumerate(zip(K_MTiles, Vt_MTiles)):
F.load_stationary(Q_STiles[i % 2])
F.load_tile(K_j, K_STiles[j % 2])
F.attn_score(K_STiles[j % 2], log_expsum)
F.load_tile(Vt_j, Vt_STiles[j % 2])
F.attn_value(Vt_STiles[j % 2], Ot_ATile)
F.reciprocal(log_expsum)
F.attn_lse_norm(Ot_ATile)
F.store_tile(Ot_ATile, Ot_MTiles[i])
# Run simulation and return result
Ot = attention(Q, K, Vt)
O = Ot.to_numpy().T # Host-side transpose
The JIT compiler processes Python functions decorated with @F.kernel. The decorator also specifies the target device (e.g., a Verilator-based simulator (Wilson, 2024)), and handles device-host communication automatically.
In the example above, the host loads input tensors into device memory using DRAMSim2 (Rosenfeld et al., 2011), dispatches the compiled instructions to the FSA device, and retrieves the output tensors upon completion. Because FSA does not currently support hardware matrix transposition, the input matrix $V$ is transposed in advance, and the output matrix $O$ is converted to NumPy (Harris et al., 2020) arrays for post-processing. On commercial hardware, such transpositions can be performed by the DMA engine during memory transfers (AWS Neuron Team, 2025), and therefore should not incur significant overhead.
6. Evaluation
6.1. FlashAttention Performance
We compare the FlashAttention performance of FSA with state-of-the-art commercial systolic-array-based accelerators, including Google’s TPUv5e (Google Cloud, 2025) and AWS NeuronCore-v2 (AWS Neuron Documentation, 2024a). We select TPUv5e because it has the same systolic array size, $128\times 128$, as AWS NeuronCore-v2. For fairness, we also configure FSA with a $128\times 128$ systolic array. All accelerators use 16-bit floating point activation and 32-bit accumulation.
The accelerator configurations used for evaluation are summarized in Table 1. We evaluate the FlashAttention forward pass performance of a single attention head on all three accelerators, using a head dimension of 128, without causal masking, and with sequence lengths ranging from 2048 to 16384.
To clarify, we use the term FLOPs to denote the number of floating-point operations, and FLOPs/s to indicate the number of floating-point operations performed per second. Although all accelerators share the same systolic array size, their absolute performance still varies due to differences in clock frequency, which in turn depend on the specific node technology employed. Therefore, we simulate FSA at 1.5 GHz, a target that is achievable in 16 nm technology, as discussed in subsection 6.3, and compare matrix multiplication FLOPs/s utilization rather than wall-clock execution time. This metric reflects the end-to-end absolute performance of each accelerator under the same clock frequency.
Table 1. Accelerator configurations used for evaluation.
Accelerator TPUv5e Neuron-v2 FSA
Systolic array size 128128 128128 128128
Number of arrays 4 1 1
TFLOPs/s (MAC only) 196.6 91.75 32.77
Frequency 1.5GHz1 2.8GHz 1.5GHz
Memory Bandwidth 819GB/s 820GB/s 820GB/s
Scratchpad SRAM 48MiB 24MiB 192KiB2
Accumulation SRAM 2MiB 64KiB
Require Vector Unit? Yes Yes No
  • 1
    The clock frequency of TPUv5e is inferred from its reported performance using the formula: $\text{freq}_{\text{GHz}}=\frac{\text{TFLOPS}\cdot 10^{12}}{2\cdot 128\cdot 128\cdot 10^{9}}$.
  • 2
    Since FSA operates on $128\times 128$ tiles and only targets FlashAttention operations in this experiment, 192 KiB of SRAM is sufficient to support the algorithm using double buffering.
Figure 11. FlashAttention FLOPs/s utilization of FSA compared to TPUv5e and AWS NeuronCore-v2.
[…]
The line plot shows the FLOPs/s utilization of FSA compared to TPUv5e and AWS NeuronCore-v2.
The number of floating-point operations in FlashAttention (Dao, 2023) is given by:
$\text{Total FLOPs}=4\times\text{SeqLen}^{2}\times d$
where SeqLen is the sequence length and $d$ is the head dimension.
We calculate the achieved FLOPs/s on each accelerator as follows:
$\text{Achieved FLOPs/s{}}=\frac{\text{Total FLOPs}}{\text{Execution Time}}$
and the FLOPs/s utilization as:
$\text{FLOPs/s{} Utilization}=\frac{\text{Achieved FLOPs/s{}}}{\text{Peak FLOPs/s{}}}$
We use the official FlashAttention kernel implementations for the two commercial accelerators, which are optimized for their respective architectures:
  • jax.experimental.pallas.ops.tpu.flash_attention for the TPUv5e (JAX Developers, [n. d.]).
  • neuronxcc.nki.kernels.attention.flash_fwd for the AWS NeuronCore-v2 (AWS Neuron Team, 2025).
The experimental results are shown in Figure 11. Benefiting from intensively overlapped computations, FSA achieves average FLOPs/s utilization that is 1.77 times and 4.83 times higher compared to TPUv5e and AWS NeuronCore-v2, respectively. Moreover, the results confirm that our RTL implementation of FSA closely aligns with the theoretical performance outlined in subsection 3.5, validating the correctness of our RTL design.
6.2. Accuracy Analysis
The SystolicAttention produces mathematically equivalent results to standard attention and preserves the order of floating-point operations. However, due to the piecewise linear interpolation used for the exp2 computation, the final results may differ slightly from those generated by commercial accelerators.
In this section, we first analyze the accuracy of the piecewise linear approximation of the single exp2 function. We then assess its overall impact on the final attention results.
6.2.1. Error Analysis of Piecewise Linear Approximation
Since the input to the exp2 function in the FlashAttention is always negative, we exhaustively evaluate the piecewise linear approximation over all negative normal fp16 values. We exclude subnormal values, as they are typically flushed to zero in most accelerators (Google Cloud, 2019).
Figure 12. exp2 piecewise linear interpolation error analysis.
[…]
The plot shows the mean absolute error (MAE) and mean relative error (MRE) of the piecewise linear approximation of the softmax function as a function of the number of pieces.
Figure 12 shows the mean absolute error (MAE) and mean relative error (MRE) of the piecewise linear approximation of the exp2 function, plotted against the number of segments used in the approximation. The MAE decreases significantly as the number of segments increases, while the MRE remains relatively stable. This occurs because the breakpoints in the approximation are uniformly spaced, whereas the fp16 representation is logarithmic in nature, making relative error more sensitive for values close to zero. We choose to use 8 segments in our FSA implementation, which achieves a MAE of 0.00014 and an MRE of 0.02728. Although non-uniformly spaced breakpoints could further reduce the MRE, we leave this exploration for future work.
6.2.2. Overall FlashAttention Accuracy
To evaluate the accuracy of the FlashAttention attention on FSA, we compare it against torch.nn.functional.scaled_dot_product_attention in PyTorch (Ansel et al., 2024). We use the same head dimension and sequence lengths as in the performance evaluation. The input matrices are randomly generated using the same distribution as in the FlashAttention-3 (Shah et al., 2024) accuracy evaluation:
$Q,K,V\sim\mathcal{N}(0,1)+\mathcal{N}(0,100)\cdot\mathrm{Bernoulli}(0.001).$
Table 2. Mean absolute error (MAE), root mean squared error (RMSE), and mean relative error (MRE) of the FlashAttention results on FSA.
SeqLen MAE RMSE MRE
2048 7.983e-03 1.315e-02 1.558e-02
4096 1.379e-02 2.290e-02 2.596e-02
6144 1.849e-02 3.085e-02 3.545e-02
8192 2.253e-02 3.772e-02 4.413e-02
10240 2.595e-02 4.373e-02 5.259e-02
12288 2.890e-02 4.873e-02 5.920e-02
14336 3.165e-02 5.351e-02 6.529e-02
16384 3.403e-02 5.784e-02 7.181e-02
The overall accuracy results are summarized in Table 2. The MAE lies in the range of $7\times 10^{-3}$ to $4\times 10^{-2}$, while the MRE ranges from $1\times 10^{-2}$ to $8\times 10^{-2}$. These results indicate that the piecewise linear approximation of the exp2 function has a negligible impact on the final attention output.
6.3. FSA Area Overhead
The area overhead of FSA arises from the following additional components:
  • The upward data path, in addition to the standard downward path, enabling partial sums to propagate in both directions.
  • A top row of CMP units, responsible for identifying the maximum value in each systolic array column.
  • A Split unit within each PE, which separates input values into integer and fractional parts for the exp2 operation.
To evaluate the area overhead, we synthesize the systolic array portion of FSA, excluding SRAMs and DMA, at 1.5 GHz using a 16 nm commercial process. The resulting area breakdown is shown in Table 3.
Table 3. FSA area breakdown.
Group Component Area (%) Area ($\mu\mathrm{m}^{2}$)
Standard PEs 86.81 24445044
Other logic 1.11 313457
Total 87.92 24758501
FSA additional Upward data path 6.24 1756641
Split units 5.30 1493150
CMP units 0.53 149524
Total 12.07 2029891
As shown in Table 3, the standard systolic array occupies 87.92% of the total area, while FSA’s additional components contribute 12.07%. The dominant sources of overhead are the upward data path and the Split units, which account for 6.24% and 5.30% of the total area, respectively. This is due to their replication across all PEs in the $128\times 128$ array. In contrast, the CMP units contribute only 0.53%, as they are instantiated solely in the top row of the array.
7. Related Work
Significant efforts have been made to fuse FlashAttention on hardware accelerators. COSA (Wang et al., 2023) and COSA Plus (Wang et al., 2025) fuse FlashAttention across two cooperative systolic arrays, but data from the first array must still pass through a special function unit (SFU) for softmax. Matching SFU throughput to large systolic arrays can increase hardware cost. FuseMax (Nayak et al., 2024) builds on the cascade Einsum abstraction introduced in TeAAL (Nayak et al., 2023) to describe the FlashAttention algorithm and map it onto spatial accelerator models. FuseMax uses a different dataflow from the one used in our paper: by employing an output-stationary dataflow, it overlaps four FlashAttention iterations on the systolic array to obtain very high utilization. The overlapping of iterations necessitates storing multiple input tiles in SRAM and maintaining intermediate results within the array, leading to higher storage overhead and control complexity. Additionally, FuseMax assumes all FlashAttention operations are performed in fp16 format. In contrast, the original FlashAttention algorithm uses fp16 for matrix multiplications but fp32 for accumulations. Reconciling this difference may require intermediate results to be stored in fp32, which can further increase register usage. ExpMul (Alexandridis et al., 2025) fuses the exp and the second matrix multiplication in FlashAttention, which is orthogonal to our work and could be integrated into FSA as well.
8. Discussion
8.1. Necessity of a Vector Unit
While FSA is capable of executing FlashAttention entirely on the systolic array without relying on a vector unit, a vector unit or SFU is still necessary in a full accelerator system to support activation functions outside of FlashAttention. However, because FSA offloads all attention-related computation from the vector unit, the overall vector FLOPs/s requirement may be significantly reduced.
8.2. FSA Variant: Area-Optimized Design
This paper focuses on the most performant FSA variant, which uses both upward and downward data flow to maximize throughput.
However, during development, we also explored a simpler variant with a single data flow direction, similar to a standard systolic array. This configuration still supports the SystolicAttention algorithm, but introduces additional latency: the second matrix multiplication ($O=PV$) cannot begin immediately after the first elements of $P$ are computed and must wait for the entire $P$ matrix.
This results in the inner loop latency increasing from $5N+10$ to $6N+10$ cycles. Despite this, this area-optimized design would still deliver higher FLOPs/s utilization than commercial accelerators.
8.3. Limitations of FSA
Although FSA delivers high FLOPs/s utilization for FlashAttention, it is not well-suited for the decoding phase of LLM inference.
FSA is ideal for the (1) LLM training and (2) prefill phase of LLM inference, where the query matrix $Q$ has a long sequence length, making the attention computation compute-intensive (Patel et al., 2024). In this case, the high FLOPs/s throughput of FSA effectively accelerates the workload.
In contrast, during the decoding phase, the query $Q$ has a sequence length of 1, making the computation memory-bound. Moreover, since FSA operates on $128\times 128$ tiles, $Q$ must be padded to fit this shape, resulting in substantial wasted computation and inefficient resource usage.
9. Conclusion
We have presented FSA, an enhanced systolic array architecture that enables complete FlashAttention computation within a single systolic array. At the core of our approach is the SystolicAttention algorithm, which achieves extensive element-wise overlap of all FlashAttention operations through careful scheduling and static data flow management.
The experimental results validate our approach, showing significant improvements in FLOPs/s utilization compared to state-of-the-art commercial accelerators while maintaining high accuracy and requiring minimal on-chip memory.